OpenStudy (anonymous):

17. A building lot in a city is shaped as a 30° -60° -90° triangle. The side opposite the 30° angle measures 41 feet.   a. Find the length of the side of the lot opposite the 60° angle.   b. Find the length of the hypotenuse of the triangular lot.   c. Find the sine, cosine, and tangent of the 30° angle in the lot. Write your answers as decimals rounded to four decimal places.

OpenStudy (anonymous):

\[\large \sf Side~opposite~60~degrees=Small~side \times \sqrt{3}\] \[\large \sf Hypotenuse = Small~side \times 2\]  \[\large \sf Sine=\frac{opp}{hyp}\]\[\large \sf Cosine=\frac{adj}{hyp}\]\[\large \sf tan=\frac{opp}{adj}\]

OpenStudy (anonymous):

All that's left is for you to plug in values and crunch the numbers